the length of a rectangle is 3 less than twice the width the perimeter is 30cm what is the length and width?

Answer by ReadingBoosters(3246)

p = 2(l+w) = 2l+2w = 30
l = 2w - 3
Substitute l into the perimeter equation
2(2w-3)+2w = 30
Distribute the 2
4w - 6 + 2w = 30
Combine like terms
6w = 30 + 6
6w=36
6w/6 = 36/6
w = 6
l = 2(6) - = 9